# PromptReps: Prompting Large Language Models to Generate Dense and Sparse Representations for Zero-Shot Document Retrieval

### v2

@@ 5.1 Zero-shot retrieval effectiveness on BEIR
We present our results on BEIR in Table 1. The first observation highlights that BM25 is a very strong zero-shot retrieval method, capable of outperforming LLM2Vec, based on the Llama3-8B-Instruct LLM, across numerous datasets, achieving a higher average nDCG@10 score. This outcome implies that even with a large-size LLM, bi-directional attention enabled, additional pre-training, and SimCSE-based unsupervised contrastive training, there remains a gap in transforming a decoder-only LLM into an effective retrieval method.
On the other hand, E5-PTlarge, based on the BERT-large model, is the first method that can outperform BM25 without any supervised training data. However, it has been trained on a massive, carefully mined text pair dataset with a large batch size, which may require more data-collecting efforts and computational resources than LLM2Vec.
Our proposed PromptReps with Llama3-8B-Instruct LLM has lower nDCG@10 scores when only using dense or sparse retrieval. However, the hybrid system (combining dense and sparse) contributes notable retrieval effectiveness improvements, surpassing BM25 and approaching the state-of-the-art E5-PTlarge. Notably, this is achieved without any form of extra training but solely relying on prompts.
Figure 2: MRR@10 scores on MS MARCO of PromptReps with different LLMs.
The scaling law of LLM Kaplan et al. (2020) also applies here. When changing Llama3-8B-Instruction to Llama3-70B-Instruction, the dense and sparse retrieval effectiveness of our PromptReps further improves, with the hybrid approach surpassing E5-PTlarge.
We further note that when interpolating dense, sparse and BM25, the average nDCG@10 achieved a remarkable score of $50.6050.60$. These results demonstrate that it is possible to build a strong retrieval system with LLMs and BM25 without the need for any unsupervised or supervised training.
@@ 5.2 Sensitivity to different prompts
In the previous experiments, we always use the prompt illustrated in Figure 1. In this section, we study how different prompts impact the retrieval effectiveness. Particularly, we design six different prompts [^10] The prompt in Figure 1 is the prompt number 6 in Table 2., listed in Table 2, and conduct experiments on TREC deep learning 2019 and 2020 datasets, and MS MARCO passage retrieval dev sub-dataset. We use Llama3-8B-Instruction as the base LLM for PromptReps. The results are listed in Table 3. We also report results of Recall@1000 and other base LLMs in Appendix A.
@@
PrompReps Llama3-8B-Instruct (ours)
@@
Interestingly, our results also show that LLMs have instruction-following ability in this representation generation task. For instance, comparing prompts #1 and #2, the only difference is the phrase “in a retrieval task”, and the prompt with this phrase yields higher retrieval effectiveness across all datasets. Additionally, comparing prompts #1 and #6, the difference is the phrase “Make sure your word is in lowercase”, which matches our sparse exact matching method where we first lowercase the input text. This phrase can further improve the retrieval effectiveness. Finally, using the adjective phrase “most important” in the prompt does not significantly impact the results.
@@
5.3 Impact of different LLMs
In this section, we explore how different base LLMs impact PromptReps. For this study, we investigate five state-of-the-art open-sourced decoder-only LLMs, covering different model sizes and models with or without instruction tuning. We use prompt #6 for all LLMs [^11] Only the model specific conversational special tokens are changed. and report MRR@10 scores on the MS MARCO datasets. The results are illustrated in Figure 2; more detailed results including on TREC deep learning datasets are reported in Appendix A.
The results show that the hybrid retrieval effectiveness of our PromptReps consistently outperforms BM25, regardless of which LLM is used, with the only exception of Mistral-7B-Instruct. When using the Mistral-7B-Instruct LLM, the dense-only retriever performs poorly. Surprisingly, implementing PromptReps with Phi-3-mini-4k-instruct achieved much higher retrieval effectiveness than that of Mistral-7B-Instruct, despite having far less parameters (3.8B).
Meta-Llama-3 models are generally very effective for our method. For 8B models, the instruction-tuned model performs significantly better than the pretrained-only model, indicating that the instruction fine-tuning is helpful to further improve our method. The 70B instruction-tuned model achieved the best hybrid retrieval results, but the dense-only and sparse-only retrieval effectiveness are similar to the 8B instruction-tuned model. These results agree with the BEIR results presented in Table 1.
@@ 5.3 Impact of different LLMs
6 Alternative representations and scoring
In the previous sections, we only considered using the representations (dense and sparse) yielded from the last token in the prompt for document retrieval. These representations, in the context of generative LLMs, are responsible for predicting the first generated token. We define this setting as First-token single-representation. We have demonstrated that this simple way of generating representations is effective for document retrieval; however, these representations might be sub-optimal. For example, LLMs use sub-word tokenization algorithms such as SentencePiece Kudo and Richardson (2018). This tokenization might split a word into sub-words, meaning that the first generated token might just be a sub-word. Using the representation of the whole word might be a better representation than the first token representation. Additionally, previous works in multi-vector dense retrieval such as ColBERT Khattab and Zaharia (2020) demonstrated that using multiple representations could be beneficial for document retrieval. How can we use PromptReps to also generate single-word representations or multiple representations that can potentially enhance the retrieval effectiveness? In this section, we explore these alternative representations.
First-word single-representation and Multi-token single-representation. Instead of just using the representations for the first generated token, these two methods let the LLM finish the generation [^12] We simply use greedy generation. of the whole word or multiple words, controlled by the given prompt (“Use one word” or “Use three words”), as illustrated in Figure 3. The end of generation is detected by the token ‘’. We then pool all the representations of the generated tokens to form a single dense and sparse representation to index the input text. For the dense representation we use mean pooling and for the sparse representation we use max pooling. Once representations are obtained, the scoring is the same as First-token single-representation.
Multi-token multi-representation and Multi-word multi-representation. Instead of using a single representation for retrieval, these two methods prompt the LLM to generate multiple words and then index each generated representation separately. The difference between the two is that Multi-token multi-representation keeps all the token representations in the index, while Multi-word multi-representation first groups tokens into words by using space, and then creates a single representation for each word by using max pooling for sparse representations and mean pooling for dense representations. During retrieval, we follow the ColBERT scoring method where the relevance score of a document is computed by the sum of the maximum similarity of each query representation against each document representation (Figure 4).
Hybrid retrieval results are shown in Figure 5, and full dense and sparse retrieval results in Appendix B. Results show that all the explored methods are able to perform document retrieval. The First-token single-representation and Multi-token single-representation generally perform the best. However, we note that Multi-token single-representation requires more token generation steps and thus has higher query latency. The First-word single-representation performs the worst, suggesting that sub-word representations hurt the retrieval performance for single-word generation prompts. On the other hand, multi-representation methods with ColBERT scoring methods do not seem beneficial. Thus, we conclude that the simplest First-token single-representation is sufficient to represent the input text for document retrieval.
@@ 7 Conclusion
We introduced PromptReps, a simple yet effective method that prompts LLMs to generate dense and sparse representations for zero-shot document retrieval without any further unsupervised or supervised training. Our work reveals that modern LLMs are effective text encoders by themselves, and prompt engineering is sufficient to stimulate their text encoding ability.
@@
Finally, our prompting method could be seen as a simple approach to obtaining a better initialization of LLM-based embedding models, which is much more cost-effective than methods requiring further pre-training BehnamGhader et al. (2024); Li et al. (2023). All the previous contrastive pre-training with paired text data and synthetically generated data could be applied on top of our method and could potentially yield improved LLM-based embedding models.
@@ 8 Limitations
Secondly, the highest effectiveness for our PromptReps is achieved in the hybrid retrieval setting. Compared to previous works which use dense representations only, the hybrid setting requires both dense and sparse retrieval, thus the extra sparse retrieval introduces extra query latency (and requires additional disk/memory space for the inverted index). However, PromptReps actually only requires a limited query latency overhead if dense and sparse retrieval are implemented in parallel. In our method, obtaining both dense and sparse representations only requires a single LLM forward inference; the only extra computation is the dot product of the dense vector with the token embeddings, which is very fast on GPU. For document search, since we heavily sparsified the sparse representation, in our experiments, our sparse retriever is much faster than BM25, and the bottleneck is the dense retriever. Since the dense and sparse search could be run in parallel and the hybrid operation is a simple linear interpolation of both rankings (very fast on CPU), the query latency of the hybrid process only depends on the dense retrieval latency, and it is thus very close to previous methods.
@@ 9 Ethical considerations
In our experiments, we use PromptReps coupled with LLMs with a large number of parameters (up to 70B in our experiments) to encode the BEIR and MS MARCO datasets, which contain millions of documents. Although no LLM training was conducted, we are aware that our experiments might still have consumed significant energy, thus contributing to CO2 emissions Scells et al. (2022) and water consumption Zuccon et al. (2023).
In addition, since we leverage LLMs in a black-box manner and LLMs’ generation might contain biases Gallegos et al. (2024), the representations generated by LLMs may be biased towards certain contents or topics. Future work could consider how to mitigate biases in PromptReps via prompt engineering.
Appendix A Full results on TREC deep learning and MS MARCO
@@
Table 4: TREC deep learning and MS MARCO performance of different prompts and LLMs.
@@
Appendix B Full results of different representation methods

### v3

@@ 5.1 Zero-shot retrieval effectiveness on BEIR
We present our results on BEIR in Table 1. The first observation highlights that BM25 is a strong zero-shot retrieval method, capable of outperforming LLM2Vec, based on the Llama3-8B-Instruct LLM, across numerous datasets, achieving a higher average nDCG@10 score. This outcome implies that even with a large-size LLM, bi-directional attention enabled, additional pre-training, and SimCSE-based unsupervised contrastive training, there remains a gap in transforming a decoder-only LLM into an effective retrieval method.
[…]
Table 2: nDCG@10 scores of unsupervised LLM-based methods hybrid with BM25 on BEIR 13 publicly available datasets. D+S+BM25 stands for PromptReps hybrid with its dense, sparse and BM25.
LLM BERT-330M Llama3-8B-I Llama3-70B-I
Dataset E5-PT+BM25 D+S+BM25 D+S+BM25
arguana 47.34 38.13 39.53
climatefever 21.92 22.95 23.34
dbpedia 43.46 40.87 41.63
fever 78.04 77.07 74.06
fiqa 42.88 34.11 35.35
hotpotqa 69.18 64.38 65.29
nfcorpus 36.61 35.20 37.64
nq 47.71 45.12 48.30
quora 88.63 86.26 86.60
scidocs 20.76 17.97 18.82
scifact 76.37 70.92 73.58
trec-covid 74.09 76.17 80.29
touche 35.01 29.13 34.15
avg 52.46 49.10 50.66
5.2 Further hybrid with BM25
@@
[…]
@@ 5.3 Sensitivity to different prompts
In the previous experiments, we always use the prompt illustrated in Figure 1. In this section, we study how different prompts impact the retrieval effectiveness. Particularly, we design six different prompts, [^10] The prompt in Figure 1 is prompt #6 in Table 3. listed in Table 3, and conduct experiments on TREC deep learning 2019 and 2020 datasets, and MSMARCO passage retrieval dev sub-dataset. We use Llama3-8B-Instruction as the base LLM for PromptReps. The results are listed in Table 4. We also report results of Recall@1000 and other base LLMs in Appendix C.
@@
PromptReps Llama3-8B-Instruct (ours)
@@
Interestingly, our results also show that LLMs have instruction-following ability in this representation generation task. For instance, comparing prompts #1 and #2, the only difference is the phrase “in a retrieval task”, and the prompt with this phrase yields higher retrieval effectiveness across all datasets. Additionally, comparing prompts #1 and #6, the difference is the phrase “Make sure your word is in lowercase”, which matches our sparse exact matching mechanism where we first lowercase the input text. This phrase can further improve the retrieval effectiveness. Finally, using the adjective phrase “most important” in the prompt does not significantly impact the results.
5.4 Impact of different LLMs
In this section, we explore how different base LLMs impact PromptReps. For this study, we investigate five state-of-the-art open-sourced decoder-only LLMs, covering different model sizes and models with or without instruction tuning. We use prompt #6 for all LLMs [^11] Only model specific conversational special tokens are changed. and report MRR@10 scores on the MSMARCO datasets. The results are illustrated in Figure 2; more detailed results including on TREC deep learning datasets are reported in Appendix C.
The results show that the hybrid retrieval effectiveness of PromptReps consistently outperforms BM25, regardless of which LLM is used, with the only exception of Mistral-7B-Instruct. When using the Mistral-7B-Instruct LLM, the dense-only retriever performs poorly. Surprisingly, implementing PromptReps with Phi-3-mini-4k-instruct achieved much higher retrieval effectiveness than that of Mistral-7B-Instruct, despite having far fewer parameters (3.8B).
Meta-Llama-3 models are generally very effective for our method. For 8B models, the instruction-tuned model performs significantly better than the pretrained-only model, indicating that the instruction fine-tuning is helpful to further improve our method. The 70B instruction-tuned model achieved the best hybrid retrieval results, but the dense-only and sparse-only retrieval effectiveness is similar to the 8B instruction-tuned model. These results agree with the BEIR results presented in Table 1.
Figure 2: MRR@10 scores on MSMARCO of PromptReps with different LLMs.
6 Supervised Results
We have demonstrated the strong zero-shot effectiveness of PromptReps. Now we explore the question: Can PromptReps serve as a better initialization for LLM-based embedding models in downstream contrastive supervised fine-tuning?
To address this question, we conduct supervised fine-tuning experiments using MSMARCO. Specifically, we follow the RepLlama training recipe Ma et al. (2024) to fine-tune the LLama-3-8B-Instruct base model with InfoNCE loss and hard negative passages mined by a BM25 and dense retrieval hybrid system. The detailed training hyper-parameters are listed in Appendix D. For the RepLlama baseline, we adhere to the original implementation, which appends the prefixes “Query: ” and “Passage: ” to the query and document text, respectively, and adds the end-of-sentence token at the end of the text. The output embedding of this token is then used to represent the text. For PromptReps, we use our proposed prompt (#6 in Table 3) and the last token hidden states and logits as the dense and sparse representation of the text. For fine-tuning PromptReps, we explore two settings, PromptReps-dense only, which only uses the dense representation of PromptReps to calculate document scores during training and inference. This setting ensures a fair comparison with RepLlama, as the only difference is the prompt used. The other setting involves using both dense and sparse document scores to calculate the loss, simply adding the two losses as the final loss. During inference, we report the dense, sparse, and hybrid retrieval effectiveness separately for this setting.
We train both RepLlama and PromptReps for 1 epoch using the full MSMARCO training data, which contains 490k training examples. In addition to the full training, we also explore a low-resource training setting, where we sample 1k examples from the entire MSMARCO dataset. We then split the 1k examples into a training and a validation set with a 9:1 ratio. We monitor the validation loss after each training epoch and stop the training, selecting the best checkpoint if no lower validation loss is observed for three consecutive epochs.
The results are shown in Table 5. Surprisingly, with only 1k training examples, ranking effectiveness of RepLlama improved from 0 to a competitive score. This finding suggests that it is possible to convert an LLM into an effective embedding model with little training data. On the other hand, PromptReps-dense only achieved the best MRR@10 score on MSMARCO dev. However, the hybrid training loss coupled with hybrid retrieval achieved the highest effectiveness across different training settings on TREC DL; the only exception being the full-data setting on MSMARCO dev. These results demonstrate that PromptReps could be seen as a simple approach to obtaining a better initialization of LLM-based embedding models, which is more cost-effective than methods requiring further pre-training BehnamGhader et al. (2024); Li et al. (2023).
Table 5: Supervised fine-tuning results
zero-shot 1k data full data (490k)
MSMARCO dev MRR@10
RepLlama3 0.0 27.88 42.77
PromptReps-dense only 17.54 28.48 42.58
PromptReps-dense 17.54 25.45 41.86
PromptReps-sparse 20.85 21.55 34.15
PromptReps-hybrid 24.62 28.18 42.48
DL2019 NDCG@10
RepLlama3 0.0 63.91 73.19
PromptReps-dense only 50.62 62.63 73.50
PromptReps-dense 50.62 64.48 74.10
PromptReps-sparse 42.25 47.23 60.39
PromptReps-hybrid 55.58 65.23 74.49
DL2020 NDCG@10
RepLlama3 0.0 63.10 73.35
PromptReps-dense only 43.81 61.46 73.00
PromptReps-dense 43.81 61.04 73.65
PromptReps-sparse 45.60 50.15 62.81
PromptReps-hybrid 56.66 64.01 73.87
7 Alternative Representation and Scoring
In the previous sections, we only considered using the representations (dense and sparse) yielded from the last token in the prompt for document retrieval. These representations, in the context of generative LLMs, are responsible for predicting the first generated token. We define this setting as First-token single-representation (FTSR). We have demonstrated that this simple way of generating representations is effective for document retrieval; however, these representations might be sub-optimal. For example, LLMs use sub-word tokenization algorithms such as SentencePiece Kudo and Richardson (2018). This tokenization might split a word into sub-words, meaning that the first generated token might just be a sub-word. Using the representation of the whole word might be a better representation than the first token representation. Additionally, previous works in multi-vector dense retrieval such as ColBERT Khattab and Zaharia (2020) demonstrated that using multiple representations could be beneficial for document retrieval. How can we use PromptReps to also generate single-word representations or multiple representations that can potentially enhance the retrieval effectiveness? In this section, we explore these alternative representations.
@@ 7 Alternative Representation and Scoring
First-word single-representation (FWSR) and Multi-token single-representation (MTSR). Instead of using the representations of the first generated token, these two methods let the LLM finish the generation [^12] We simply use greedy generation. of the whole word or multiple words, controlled by the given prompt (“Use one word” or “Use three words”), as illustrated in Figure 4. The end of generation is detected by the token ‘ ’. We then pool all the representations of the generated tokens to form a single dense and sparse representation for the input text. For the dense representation we use mean pooling and for the sparse representation we use max pooling. Once representations are obtained, the scoring is the same as FTSR.
Multi-token multi-representation (MTMR) and Multi-word multi-representation (MWMR). Instead of using a single representation for retrieval, these two methods prompt the LLM to generate multiple words and then index each generated representation separately. The difference between the two is that MTMR keeps all the token representations in the index, while MWMR first groups tokens into words by using space, and then creates a single representation for each word by using max pooling for sparse representations and mean pooling for dense representations. During retrieval, we follow the ColBERT scoring method where the relevance score of a document is computed by the sum of the maximum similarity of each query representation against each document representation (Figure 4).
Hybrid retrieval results are shown in Figure 5, and full dense and sparse retrieval results in Appendix E. Results show that all the explored methods are able to perform document retrieval. The FTSR and MTSR generally perform the best. However, we note that MTSR requires more token generation steps and thus has higher query latency. The FWSR performs the worst, suggesting that sub-word representations hurt the retrieval performance for single-word generation prompts. On the other hand, multi-representation methods with ColBERT scoring methods do not seem beneficial. Thus, we conclude that the simplest FTSR is sufficient to represent the input text for document retrieval.
@@ 8 Conclusion
We introduced PromptReps, a simple yet effective method that prompts LLMs to generate dense and sparse representations for zero-shot document retrieval without any further training. We show that modern LLMs are effective text encoders by themselves, and prompt engineering is sufficient to stimulate their text encoding ability.
@@
Finally, our prompting method could be seen as a simple approach to obtaining a better initialization of LLM-based embedding models and all the previous contrastive pre-training with paired text data and synthetically generated data could be applied on top of our method and could potentially yield improved LLM-based embedding models.
@@ 9 Limitations
Second, the highest effectiveness for PromptReps is achieved in the hybrid retrieval setting. Compared to previous works which use dense representations only, the hybrid setting requires both dense and sparse retrieval, thus the extra sparse retrieval introduces extra query latency (and requires additional disk/memory space for the inverted index). However, PromptReps actually only requires a limited query latency overhead if dense and sparse retrieval are implemented in parallel. In our method, obtaining both dense and sparse representations only requires a single LLM forward inference; the only extra computation is the dot product of the dense vector with the token embeddings, which is very fast on GPU. For document search, since we heavily sparsified the sparse representation, in our experiments, our sparse retriever is much faster than BM25, and the bottleneck is the dense retriever. Since the dense and sparse search could be run in parallel and the hybrid operation is a simple linear interpolation of both rankings (very fast on CPU), the query latency of the hybrid process only depends on the dense retrieval latency, and it is thus very close to previous methods.
@@ 10 Ethical Considerations
In our experiments, we use PromptReps coupled with LLMs with a large number of parameters (up to 70B in our experiments) to encode the BEIR and MSMARCO datasets, which contain millions of documents. Although no LLM training was conducted, we are aware that our experiments might still have consumed significant energy, thus contributing to CO2 emissions Scells et al. (2022) and water consumption Zuccon et al. (2023). In addition, since we leverage LLMs in a black-box manner and LLMs’ generation might contain biases Gallegos et al. (2024), the representations generated by LLMs may be biased towards certain contents or topics. Future work could consider how to mitigate biases in PromptReps via prompt engineering.
Acknowledgments
The authors from the University of Waterloo acknowledge the support of the Natural Sciences and Engineering Research Council (NSERC) of Canada.
Appendix A Python code example
In Table 6 we provide a Python code implementation of PromptReps with Meta-Llama-3-8B-Instruct LLM. The example uses Huggingface transformers library (v4.40.1) with torch (v2.3.0), numpy (v1.26.4) and nltk (v3.9.1) to generate dense and sparse representations.
Table 6: Python code example of generating dense and sparse representations with PromptReps.
from transformers import AutoModelForCausalLM, AutoTokenizer import torch import numpy as np from nltk import word_tokenize from nltk.corpus import stopwords import string stopwords = set(stopwords.words(’english’) + list(string.punctuation)) model_id = ’meta-llama/Meta-Llama-3-8B-Instruct’ tokenizer = AutoTokenizer.from_pretrained(model_id) model = AutoModelForCausalLM.from_pretrained(model_id).to(’cuda’) passage = ’The quick brown fox jumps over the lazy dog.’ messages = [ {’role’: ’system’, ’content’: ’You are an AI assistant that can understand human language.’}, {’role’: ’user’, ’content’: f’Passage: "{passage}". Use one word to represent the passage’ f in a retrieval task. Make sure your word is in lowercase.’}, {’role’: ’assistant’, ’content’: ’The word is "’} ] input_ids = tokenizer.apply_chat_template( messages, add_generation_prompt=False, return_tensors=’pt’ )[:, :-1].to(’cuda’) # the last special token is removed outputs = model(input_ids=input_ids, return_dict=True, output_hidden_states=True) # generating dense representation dense_representation = outputs.hidden_states[-1][:, -1, :][0] # generating sparse representation, log and relu the values next_token_logits = torch.log(1 + torch.relu(outputs.logits))[:, -1, :][0] # lower case and stopwords removal words_in_text = [word for word in word_tokenize(passage.lower()) if word not in stopwords] # extract token ids in the given passage token_ids_in_text = set() for word in words_in_text: token_ids_in_text.update(tokenizer.encode(word, add_special_tokens=False)) token_ids_in_text = torch.tensor(list(token_ids_in_text)) # get top tokens and quantization top_k = min(len(token_ids_in_text), 128) top_k_values, top_k_indices = next_token_logits[token_ids_in_text].topk(top_k, dim=-1) values = np.rint(top_k_values.cpu().detach().float().numpy() * 100).astype(int) tokens = [tokenizer.decode(i) for i in token_ids_in_text[top_k_indices.cpu().detach().float().numpy()]] # final sparse representation print({token: value for token, value in zip(tokens, values)}) # {’fox’: 312, ’dog’: 280, ’brown’: 276, ’j’: 273, ’quick’: 265, ’lazy’: 257, ’umps’: 144}
Appendix B Impact of Hybrid weights
In Figure 6 we plot the MRR@10 scores obtained with different weights for PromptReps with Llama-8B-Instruct on the MSMARCO dev set. As the plot suggests, the best effectiveness is achieved with weight set to 0.4, noting that 0.5 (the setting used in our zero-shot experiments) actually provides a fairly good choice for hybrid retrieval.
Figure 6: Average MRR@10 scores on MSMARCO dev queries of PromptReps with different dense and sparse fusion weights.
Appendix C Full results on TREC deep learning and MSMARCO
@@
Table 7: TREC deep learning and MSMARCO performance of different prompts and LLMs. +BM25 is the system that hybrid dense, sparse, and BM25.
@@
Appendix D Fine-tuning hyper-parameters
In Table 8 we report the fine-tuning hyper-parameters we used for both RepLLama and PromptReps in Section 6. We use the training data with hard negatives provided in Tevatron Huggingface hub. [^13] https://huggingface.co/datasets/Tevatron/msmarco-passage-aug
Table 8: Hyper-parameters for supervised fine-tuning on MSMARCO passage ranking dataset.
LLM LLama3-8B-Instruct
learning rate 1e-4
warmup ratio 0.1
per GPU batch size 8
# of GPUs 4
gradient accumulation steps 4
# of negative per example 15
total in batch negative 511
distance method cosine similarity
score temperature 0.01
query length 32
passage length 156
LoRA rank 8
Appendix E Full results of different representation methods